OpenStudy (anonymous):

375 students went to Six Flags over Georgia. Fifteen buses were filled and 15 students rode with their parents. How many syudents wereon each bus?

OpenStudy (anonymous):

If there are 375 students in total and 15 of those students travelled with their parents, then (375-15) = 360 students were on buses. Divide that number by the number of buses to determine how many students were on each bus.

OpenStudy (anonymous):

thanks! I need help on more problems :/

OpenStudy (anonymous):

-3(5 - 15n) = 15(3n - 2)